# Robust Anomaly Detection in Network Traffic: Evaluating Machine Learning Models on CICIDS2017

### v1

@@ I Introduction
Following the global COVID-19 pandemic, people have become increasingly dependent on network infrastructure. With the widespread adoption of smart devices and the Internet of Things (IoT), a great volume of data are generated every day. Despite the significant convenience brought about by IoT devices, their widespread adoption inevitably leads to the transmission of vast amounts of sensitive information across networks. A substantial portion of this data—including personal, financial[1], and operational records—is now stored directly on cloud platforms such as AWS, Azure, and Google Cloud. This growing reliance on cloud-based infrastructure raises a critical question: how can we ensure the security and privacy of this data as it flows through increasingly complex and distributed networks[2].
@@ II Literature Review
Together, these studies demonstrate the evolution from signature-based IDS toward hybrid and anomaly-based methods. However, the trade-offs between generalization, accuracy, and interpretability remain open challenges—particularly in the context of unseen or evolving attack patterns.
@@ III-C1 Multi-Layer Perceptron (MLP)
$\mathcal{L}_{\mathrm{BCE}}=-\left[y\log(\hat{y})+(1-y)\log(1-\hat{y})\right]$
where $y$ is the true label (00 for benign, $1$ for malicious), and $\hat{y}$ is the predicted probability.
@@
To capture local temporal and spatial patterns within the network flow features, we employ a one-dimensional convolutional neural network (1D-CNN) composed of multiple convolutional and pooling stages followed by a fully connected classifier. Concretely, each convolution layer computes
@@ V Discussion
  • MLP relies on labeled samples of both benign and known attack types during training. Its excellent performance on the Overall Test Set stems from its ability to learn discriminative patterns directly tied to these labels. However, it lacks exposure to unseen attack behaviors and therefore fails to generalize, which explains its drastic drop in recall when tested on novel attacks. Similar generalization challenges have been noted in other systems [14].
@@
  • OCSVM trains solely on benign samples, defining a boundary around normal traffic. It flags out-of-distribution patterns as anomalies. This method strikes an optimal balance: it avoids overfitting to known attacks while maintaining precise decision boundaries, resulting in robust detection across both known and novel attack scenarios [7].
Conclusions from Model Behavior: These observations suggest that anomaly detection models defining decision boundaries based only on benign data can be more robust to unseen attacks than supervised models relying on known attack patterns. Among the supervised methods, both MLP and CNN achieve strong performance on familiar threats but suffer significant drops in recall on novel attacks. Within the unsupervised paradigm, our results show that boundary-based OCSVM consistently outperforms density-based LOF for IDS tasks: OCSVM’s global decision surface yields fewer false alarms in mixed-density traffic and more stable recall on novel attacks, making it especially well-suited for real-world intrusion detection.
Future Work: Future research could explore hybrid architectures that combine supervised and unsupervised components, leveraging the sequence modeling capability of frameworks like SETransformer [15, 18]. Another promising direction involves post-training model adaptation[16], where attribute unlearning techniques [17][14]allow intrusion detectors to incorporate new patterns without retraining from scratch.

### v2

@@ I Introduction
Following the global COVID-19 pandemic, people have become increasingly dependent on network infrastructure. With the widespread adoption of smart devices and the Internet of Things (IoT), a great volume of data are generated every day. Despite the significant convenience brought about by IoT devices, their widespread adoption inevitably leads to the transmission of vast amounts of sensitive information across networks. A substantial portion of this data—including personal, financial[1], and operational records—is now stored directly on cloud platforms such as AWS, Azure, and Google Cloud. This growing reliance on cloud-based infrastructure raises a critical question: how can we ensure the security and privacy of this data as it flows through increasingly complex and distributed networks[2, 26].
@@ II Literature Review
Together, these studies demonstrate the evolution from signature-based IDS toward hybrid and anomaly-based methods. However, the trade-offs between generalization, accuracy, and interpretability remain open challenges—particularly in the context of unseen or evolving attack patterns. Similar challenges have also been observed in risk modeling tasks beyond cybersecurity, such as credit risk assessment using structural graph-based models [21].
@@ III-C1 Multi-Layer Perceptron (MLP)
[…]
where $y$ is the true label (0 for benign, $1$ for malicious), and $\hat{y}$ is the predicted probability.
@@
benign tr local temporal and spatial patterns within the network flow features, we employ a one-dimensional convolutional neural network (1D-CNN) composed of multiple convolutional and pooling stages [20] followed by a fully connected classifier. Concretely, each convolution layer computes
@@ V Discussion
  • MLP relies on labeled samples of both benign and known attack types during training. Its excellent performance on the Overall Test Set stems from its ability to learn discriminative patterns directly tied to these labels. However, it lacks exposure to unseen attack behaviors and therefore fails to generalize, which explains its drastic drop in recall when tested on novel attacks. Similar generalization challenges have been noted in other systems [14], including concept-level backdoor vulnerabilities in interpretable models [23].
@@
  • OCSVM trains solely on benign samples, defining a boundary around normal traffic. It flags out-of-distribution patterns as anomalies. This method strikes an optimal balance: it avoids overfitting to known attacks while maintaining precise decision boundaries, resulting in robust detection across both known and novel attack scenarios [7, 22, 19].
Conclusions from Model Behavior: These observations suggest that anomaly detection models defining decision boundaries based only on benign data can be more robust to unseen attacks than supervised models relying on known attack patterns. Among the supervised methods, both MLP and CNN achieve strong performance on familiar threats but suffer significant drops in recall on novel attacks. Within the unsupervised paradigm, our results show that boundary-based OCSVM consistently outperforms density-based LOF for IDS tasks: OCSVM’s global decision surface yields fewer false alarms in mixed-density traffic and more stable recall on novel attacks, making it especially well-suited for real-world intrusion detection, echoing broader efforts to build robust and interpretable AI systems in safety-critical domains such as autonomous driving [24].
Future Work: Future research could explore hybrid architectures that combine supervised and unsupervised components, leveraging the sequence modeling capability of frameworks like SETransformer [15, 18]. Another promising direction involves post-training model adaptation[16], where attribute unlearning techniques [17, 14, 25]allow intrusion detectors to incorporate new patterns without retraining from scratch.